OpenStudy (anonymous):

evaluate the series (1 + (2k/n))(1/n) from k=1 to n

OpenStudy (turingtest):

\[\sum_{k=1}^n(1+\frac{2k}n)(\frac1n)\]?

OpenStudy (anonymous):

yup

OpenStudy (turingtest):

n is a constant so let's treat it as such, meaning we can pull it out of the series when multiplying\[\sum_{k=1}^n(1+\frac{2k}n)(\frac1n)=\frac1{n^2}\sum_{k=1}^n(n+2k)\]so far so good?

OpenStudy (anonymous):

yup

OpenStudy (turingtest):

now any idea how we go about handling what's left over? remember that n is still a constant, while k will be changing...

OpenStudy (anonymous):

we never learnt this in the first place so ....

OpenStudy (turingtest):

consider the series\[\sum_{k=1}^n5\]what is the sum of that?

OpenStudy (anonymous):

5?

OpenStudy (turingtest):

no, how about\[\sum_{k=1}^35\]think about what the sum sign means

OpenStudy (turingtest):

\[\sum_{k=1}^35=5+5+5=3(5)\]

OpenStudy (anonymous):

15?

OpenStudy (turingtest):

right :)

OpenStudy (anonymous):

so the sigma means to do something an n number of times?

OpenStudy (turingtest):

it means to add up n number of those terms, so yeah the specific "something" that you are doing n times is adding the terms together

OpenStudy (turingtest):

so notice what happened there\[\sum_{k=1}^35=5+5+5=3(5)=15\]so what about in general?\[\sum_{k=1}^n5\]since 5 will occur n times, the sum will be...?

OpenStudy (anonymous):

5*n?

OpenStudy (turingtest):

yes :)

OpenStudy (turingtest):

now, remembering that n is a constant, what does that mean about\[\sum_{k=1}^nn=?\]

OpenStudy (anonymous):

n^2?

OpenStudy (turingtest):

yes :) very good so now let's return to our problem...

OpenStudy (turingtest):

\[\sum_{k=1}^n(1+\frac{2k}n)(\frac1n)=\frac1{n^2}\sum_{k=1}^n(n+2k)\]we can split up the next part into two remaining sums:\[\frac1{n^2}\left(\sum_{k=1}^nn+\sum_{k=1}^n2k\right)=\frac1{n^2}\left(n^2+\sum_{k=1}^n2k\right)=1+\frac1{n^2}\sum_{k=1}^n2k\]still with me?

OpenStudy (anonymous):

yup

OpenStudy (turingtest):

now what can we do about\[\sum_{k=1}^n2k\]? first notice that 2 is a constant, so what can we do with it?

OpenStudy (anonymous):

take it out

OpenStudy (turingtest):

right, so we have\[\frac1{n^2}\left(\sum_{k=1}^nn+\sum_{k=1}^n2k\right)=\frac1{n^2}\left(n^2+\sum_{k=1}^n2k\right)=1+\frac2{n^2}\sum_{k=1}^nk\]so now it all comes down to\[\sum_{k=1}^nk=1+2+3+...+(n-1)+n\]for which there is a very well-known formula (attributed to the mathematician Gauss) that says the the sum of the first n natural numbers is\[{n(n+1)\over2}\]and that is precisely what we have with our sum so use that formula for the last sum and you're ready to do the final algebra needed to solve it

OpenStudy (anonymous):

okay, i see

OpenStudy (turingtest):

you should really try to prove that theorem if/when you learn mathematical induction. It's quite useful.

OpenStudy (anonymous):

so the answer would be 1 + (n+1)/n ?

OpenStudy (turingtest):

yep :) nice job!

OpenStudy (anonymous):

hmm thats funny, our teacher got 1 + ((n+2)(n))/2

OpenStudy (turingtest):

that is funny, but instead of laughing I want a second opinion to see who's actually wrong @amistre64 am I wrong about this series?

OpenStudy (anonymous):

i was confused about where she got the +2 from. i get the logic behind yours though

OpenStudy (turingtest):

@asnaseer did I mess up?

OpenStudy (turingtest):

@Callisto

OpenStudy (asnaseer):

the answer [1 + (n+1)/n = 1 + 1 + 1/n = 2 + 1/n] is correct

OpenStudy (turingtest):

oh good, so I guess either you copied part of the problem wrong, or the teacher was wrong @halcyon98

OpenStudy (asnaseer):

and your approach is also correct

OpenStudy (turingtest):

thanks asnaseer :)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

i think she was wrong because here answers were way off on a bunch. thanks alot for clearing that up

OpenStudy (asnaseer):

I just entered this in wolfram as a double check: http://www.wolframalpha.com/input/?i=sum+%281+%2B+%282k%2Fn%29%29%281%2Fn%29+k%3D1+to+n answer still correct :)

OpenStudy (anonymous):

great, thanks!

OpenStudy (asnaseer):

yw :)